OpenStudy (anonymous):

Simplify? (-3x to the second) times (5x to the negative 3rd) PLEASE HELP

OpenStudy (anonymous):

\[(-3x ^{2}) (5x ^{-3})\]

OpenStudy (anonymous):

\[-3x ^{2}\times5x ^{-3}=-15x ^{2-3}=-15x ^{-1}\] Since it is multiplication, the constants -3 and 5 can be pulled out front and multiplied together. When multiplying a variable raised to two different powers, simply just add the powers together. A simple example is x*x = x^2 since x = x^1 and 1+1=2.

OpenStudy (anonymous):

Oh I understand. Thanks